OpenStudy (anonymous):

24 · log(2x)=60 solve for x

OpenStudy (anonymous):

24 * log(2x)?

OpenStudy (anonymous):

Llort, i already tried that site and it doesnt give me the answer i need

OpenStudy (anonymous):

\[24\log(2x)=60\] \[\log(2x)=\frac{60}{24}=\frac{5}{2}\] \[2x=10^{\frac{5}{2}}\] \[x=\frac{10^{\frac{5}{2}}}{2}\]

OpenStudy (anonymous):

\[\log_{10}(2x) = \frac{60}{24} \Rightarrow 2x = 10^{\frac{60}{24}} \Rightarrow x = \frac{1}{2}10^{\frac{60}{24}}\]

OpenStudy (anonymous):

i was too lazy to reduce fractions :(

OpenStudy (anonymous):

joemath has it

OpenStudy (anonymous):

yeah, \[\frac{60}{25}\]!

OpenStudy (anonymous):

hey hero